the length of a flag is twice the width. Find the width of a flag with an area of 1250 cm²
Answer by mananth(16949)

the length of a flag is twice the width. Find the width of a flag with an area of 1250 cm²
let the width be x
length = 2x
Area = L*W
2x^2=1250
X^2=1250/2
X^2=625
X=25
mananth@hotmail.com